OpenStudy (anonymous):

solving logarithmic and exponential equations 1. 8-3^x = -1 2. log(lowered 3)81=x 3. log x- log3=3

OpenStudy (kenljw):

8-3^x = -1 3^x =9 x log(3) =9 x=9/log(3) log x- log3=3 log(x) = 3 + log(3) x = 10^(log(3)+3) x= ]10^log(3)] 10^3 x=3 10^3

OpenStudy (anonymous):

are those the answers to 1 and 3?

OpenStudy (kenljw):

Does it look like the solutions

OpenStudy (anonymous):

Well I got 3 as the solution on number 1... 2.587 for number 7 and 300 for number 3?? are those right?

OpenStudy (kenljw):

I didn't do pb 2 because I didn't know what lowered 3 was, suspect 1/3 doesn't matter, if you solve on calculator my answer you don't get yours error somewhere, check a see how I solved them, if you understand it you should be able to do pb 2

OpenStudy (anonymous):

THANK YOU SO MUCH!